OpenStudy (anonymous):

9x+2y=-46 -x+y=-1 plz show work

OpenStudy (anonymous):

y = x - 1 9x + 2(x-1) = -46 9x + 2x - 2 = -46 11x = -44 x = -4 y = x-1 y = -5 (-4, -5)